Julie Cupples

6.Julie Cupples is in the process of evaluating her portfolio. Help her answer the following questions:

a.She is considering two bonds for addition to her portfolio. Bond A has a modified duration of 8.0. Bond B has a modified duration of 3.0. Julie expects interest rates to fall over the next three years. Based solely on this information (assume the investments have equivalent bond ratings), which bond should she purchase?

b.Julie is currently in the 22 percent marginal tax bracket. She can purchase a AAA corporate bond with a 4 percent yield to maturity. Alternatively, she can purchase a AAA municipal bond that yields 3.25 percent. Which is the better option for Julie?

c.Julie is a gambler. She thinks the value of a stock she has been tracking is going to drop dramatically over the next three months. Julie does not have enough money in her brokerage account to short the stock. What strategy can she use to bet that the price of the shares will fall without having to hold shares directly?

d.Julie is interested in buying a real estate investment trust for her portfolio. The REIT’s price is currently $95 per share. Julie uses a value investing approach when managing her portfolio—she likes bargains. If the stock currently pays a $4 dividend should she make the purchase if her required rate of return is 9 percent and the dividend is growing at 5 percent annually?
Answer #1

Part A

As Julie is expecting the interest rates to fall over the next three years , she should purchase 3 years modified duration bonds as the interest rates in bonds will be reducing in the coming years and hence her interest income will reduce . So by purchasing less duration bonds she can repurchase the bonds after the maturity at a better interest rate.

Part B

So as Julie is currently under 22 percent tax bracket , she will be paying tax on the yield accruing to her on AAA Corporate bond whereas AAA Municipal bonds are tax free.

Assuming the bond amount is 10,000

Yield on AAA Corporate bond = 4% of 10,000 = 400

Yield on AAA Municipal bond = 3.25% of 10,000 = 325

Yield forgone = 400-325 = 75

However she will be paying 22% tax on AAA corporate bond , hence tax expense = 22% of 400 = 88

Therefore purchasing of AAA Municipal bond is more beneficial for Julie eventhough she is foregoing 75 but saving 88 by not paying tax.

Part C

Julie can trade into futures or options for that stock by paying only a margin money to the broker .

If Julie thinks that the price of the stock will decrease dramtically then she should buy a 3 months put option.

She can also sell Futures of that stock till the next expiry.

But Options will be a better choice.

Part D

Dividend yield = 4/95*100 = 4.21%

If Julie wants to invest for a period of more than 2 years she should purchase the REITS as the dividends will be reinvested and more dividends will be paid to her which will be more than her required rate of return of 9% .

Also the dividend is growing at 5% annually which additionally provide more return.

Also if a company which gives constant dividends . will also give dividends even if the earnings are less.

Therefore Julie should buy the REITS at the current price and dividend.